OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) - 4.

OpenStudy (broskishelleh):

x = -1

OpenStudy (anonymous):

Thank you

OpenStudy (broskishelleh):

Can I have a fan and medal, and then you could tag me

OpenStudy (broskishelleh):

Oh and welcome to Openstudy:)

OpenStudy (broskishelleh):

Hi @mathmale

OpenStudy (mathmale):

Hate to "rain on your parade," but for one of you to do the problem and to give away the answer with no involvement on the part of the other, or to accept such an answer without going through the problem solution by oneself, both violate OpenStudy's rules. I'd appreciate it if you, Broskishelleh, would set up step-by-step instructions whereby Shaee will be able to arrive at the correct solution herself.

OpenStudy (broskishelleh):

Ah yes. 3(x + 1) = -2(x - 1) - 4 3(-1 + 1) = -2(-1 - 1) - 4

OpenStudy (broskishelleh):

The way I just put is true, meaning it can be solved. @Shaee Try solving 3(-1 + 1) = -2(-1 - 1) - 4

OpenStudy (anonymous):

Oh ok, I don't know how to give any fan medals

OpenStudy (anonymous):

@Broskishelleh

OpenStudy (broskishelleh):

Just click BEST RESPONSE

OpenStudy (broskishelleh):

@Shaee did you solve it though

OpenStudy (anonymous):

Yes I did, thank you again

OpenStudy (mathmale):

Happy you are happy, @Shaee, but the teacher in me is going to ask you next time to share all of your work! @Broskishelleh: thank you!

OpenStudy (broskishelleh):

@mathmale why thank you, I thought you were mad at me and Shaee

OpenStudy (broskishelleh):

@mathmale I give you a medal!

OpenStudy (mathmale):

Thank you, B!

OpenStudy (broskishelleh):

@mathmale were you mad?

OpenStudy (mathmale):

Naw, just a little annoyed. But I expect to survive. :)

OpenStudy (broskishelleh):

@mathmale can you become my fan, or no....

OpenStudy (broskishelleh):

@mathmale a medal, from you?!!!!